if the line passes through points (0,5) and (-2,9), the equation for the line is?
Answer by rfer(12644)

m=(9-5)/(-2-0)
m=4/-2
m=-2
-------------
y=mx+b
9=-2*-2+b
9=4+b
b=9/4
b=2.25
---------------
y=-2x+2.25